OpenStudy (anonymous):

Show that: x^2 + x + 1 = y ^2 has no positive integer solutions

OpenStudy (anonymous):

If x^2+x+1 takes on an perfect square value for any positive integer x, then there will be a postive integer solution. However, x^2+x+1 = (x+1)^2-x. The difference between two x^2 and (x+1)^2 is 2x+1, so (x+1)^2-x will be between x^2 and (x+1)^2 and thus can't be a perfect square.

OpenStudy (anonymous):

Isn't "The difference between two x^2 and (x+1)^2 is 2x+1" 2x^2 - x^2 + 2x + 1 = x^2 + 2x + 1

OpenStudy (anonymous):

Whoops, I don't know how that "two" got in there - I meant just x^2.

OpenStudy (anonymous):

hm, an alternate solution calls for the following assume the equation is true. Then, x<y. By manipulation, we get x^2-y^2 = 1 + y or (x-y)(x+y) = 1 + y Since y>x≥1 Here's where it's unclear for me: we get x-y > 1 and x + y > 2 + y

OpenStudy (anonymous):

Then you would prove that 1*(2+y) becomes a contradiction with the right hand side of the given equation

OpenStudy (anonymous):

and also why isn't y considered?